# Directional Textual Inversion for Personalized Text-to-Image Generation

### v1

@@ 1 Introduction
Personalization in text-to-image generation involves the targeted adaptation of models to learn representations of novel, user-provided concepts. This process allows for the creation of customized images that faithfully render specific concepts, such as unique individuals, objects, or artistic styles, in new contexts.
Current personalization approaches fall into two paradigms: parameter fine-tuning and embedding optimization. Parameter fine-tuning methods, exemplified by DreamBooth (ruiz_dreambooth_2023), optimize entire models using a few user-provided images. While effective, these approaches are computationally expensive and require significant storage per concept. In contrast, embedding optimization methods, such as Textual Inversion (gal_image_2023), offer a more efficient alternative by optimizing only token embeddings. This approach provides substantial advantages: minimal storage per concept and seamless workflow integration. These advantages have made TI a foundational component in numerous personalization frameworks (hao_vico_2023; kumari_multi-concept_2023; tewel_key-locked_2023; lee_direct_2024) and align with a broader paradigm shared with other domains, such as LLM (lester2021power) and VLM (alaluf2024myvlm).
Despite its utility, TI suffers from critical limitations. The fundamental challenge stems from the constraint of optimizing a single embedding vector to encapsulate complex visual concepts. This limitation leads to two key problems. First, TI struggles to maintain high fidelity to complex prompts, compromising its controllability and expressive range. Second, the extensive fine-tuning duration required for each concept hinders its practical applicability. Recent works (voynov_p_2023; alaluf2023neural) have attempted to address these limitations through enriched embedding spaces, but introduce significant computational overhead that undermines TI’s efficiency advantage. Moreover, these methods do not directly address the underlying optimization dynamics of TI, leaving the fundamental factors that govern semantic alignment in embedding-based personalization unclear.
This paper presents a systematic analysis of the optimization dynamics in TI, with a specific focus on the characteristics of the token embedding space. Our investigation reveals that semantic information is predominantly encoded in the direction of the embedding vectors. Furthermore, we demonstrate both theoretically and empirically that the magnitude of these embeddings is a primary source of instability; specifically, excessively high embedding norms emerge during optimization and act as a critical factor impairing image-text alignment.
@@ 1 U+2069 apples
As shown in Table 2.1, an embedding’s nearest neighbors are semantically coherent when measured by cosine similarity but not by Euclidean distance. For the token ‘apple’, its cosine-based neighbors include ‘apples’, ‘fruit’, and ‘pear’, while its Euclidean-based neighbors are often unrelated tokens with a similar magnitude. This indicates that an embedding’s direction is the primary carrier of semantic information. More results are provided in Appendix A.
@@ 2.2 Why large magnitudes lead to low text fidelity
Effect II: Residual updates stagnate (see Lemma 2).  The residual updates, $F_{\ell}(\operatorname{Norm}({\bm{x}}^{(\ell)}))$, are computed from a normalized inputs and thus have a bounded magnitude. When this bounded update is added through the skip connection to a large vector ${\bm{x}}^{(l)}$, the relative change (i.e., turning angle of the hidden state’s direction) becomes tiny, decreasing in proportion to $1/\|{\bm{x}}^{(l)}\|$. In other words, large-norm hidden states become stuck in their direction and are difficult for subsequent layers to refine. This residual stagnation accumulates across layers, severely limiting the total directional change the initial token can undergo, as formalized in the following proposition and corollary.
@@
$\angle(\alpha{\bm{x}}^{(0)},{\bm{x}}^{(L)}(\alpha))\;\leq\;\frac{\pi}{2}\,\frac{S_{L}}{\alpha\left\lVert{\bm{x}}^{(0)}\right\rVert-S_{L}}0\;\xrightarrow[\alpha\to\infty]{}\;0,$
@@
Together, these two effects explain why TI struggles with text fidelity. As token’s magnitude increases, its ability to integrate contextual information from the prompt diminishes. The personalized token becomes too dominant that it overshadows other critical details, such as stylistic elements, background context, or additional subjects, from the generated output. To this end, this analysis highlights the need for a method that explicitly controls the magnitude of personalized tokens, which we introduce in the next section.
@@ 2.3 Empirical Validation
We evaluate whether increasing embedding magnitude makes positional information unrecoverable after the first pre-norm normalization ($\mathrm{LN}$). To validate this, we train a 2-layer MLP classifier on the frozen base text encoder to predict a token’s absolute position from the output of $\mathrm{LN}(\mathbf{e}+\mathbf{p})$, where $\mathbf{e}$ and $\mathbf{p}$ each denotes token and positional embeddings. On unmodified inputs (‘Normal‘ in Figure 2), the classifier achieves $100\%$ accuracy, confirming that $\mathrm{LN}$ preserves positional information. We then scale the norm of a single token embedding by a factor $m\in\{0.5,1,2,4,8,16\}$ before applying $\mathrm{LN}$. Accuracy deteriorates rapidly once $m$ exceeds the natural scale of the encoder. Furthermore, we evaluated the classifier on TI-trained and DTI-trained personalized embeddings. TI embeddings, which has excessively large norms, collapse to near-zero positional accuracy, while DTI embeddings remain fully recoverable.
@@ 3 Method: Directional Textual Inversion
Based on our observation and analysis on previous section that token embeddings exhibit strong directional characteristics, we introduce Directional Textual Inversion (DTI), a framework that optimizes an embedding’s direction with in-distribution norm to enhance text fidelity in personalized text-to-image generation.
@@ 3.1 Optimizing only direction on the hypersphere
Here, $\mathbb{S}^{d-1}=\{{\bm{u}}\in R^{d}:\|{\bm{u}}\|_{2}=1\}$ denotes the unit sphere. We fix the magnitude $m^{\star}$ and optimize only the direction ($\bm{v}$}). Specifically, we set $m^{\star}$ to be an in-distribution magnitude derived from the frozen vocabulary of text encoder (e.g., the average norm). In this way, optimization focuses on semantic information in direction while avoiding out-of-distribution (OOD) norms.
Since the parameter space is the unit sphere, Euclidean updates drift off-manifold, making AdamW (loshchilov2017decoupled) (default optimizer used in TI-like methods) not suitable. To solve this, we use Riemannian stochastic gradient descent (RSGD) (bonnabel2013rsgd) with tangent-space projection and retraction:
@@
Here, ${\bm{g}}_{\text{euc}}$ is a Euclidean space gradient, ${\bm{g}}\in T_{{\bm{v}}_{k}}\mathcal{S}^{d-1}$ is a tangent-space gradient, and $\eta>0$ is a learning rate. In practice, we scaled the gradient $\bm{g}$} using its own norm similarly. This was inspired by Euclidean space optimizers (hinton2012rmsprop; kingma_adam_2015; loshchilov_decoupled_2019), which normalizes the gradient based on moving average of squared gradients. See Algorithm 1 and Appendix C.1 for further details.
@@ 3.2 Maximum A Posteriori formulation with a directional vMF prior
To incorporate directional prior, we formulate the optimization for the optimal direction $\bm{v}$$^{*}$ as a Maximum A Posteriori (MAP) estimation problem. Given a dataset of images $\mathcal{D}=\{{\bm{z}}_{1},\dots,{\bm{z}}_{n}\}$, the MAP estimate is found by maximizing the posterior probability:
${\bm{v}}^{*}=\arg\max_{\mathbf{v}}p(\mathbf{v}\mid\mathcal{D})\propto\arg\max_{\mathbf{v}}\left[\log p(\mathcal{D}\mid\mathbf{v})+\log p(\mathbf{v})\right].$ (3)
Minimizing the negative log-posterior is equivalent to minimizing a loss function composed of a data term and a prior term, $\mathcal{L}({\bm{v}})=\mathcal{L}_{\text{data}}({\bm{v}})+\mathcal{L}_{\text{prior}}({\bm{v}})$.
@@
$\mathcal{L}_{\text{data}}({\bm{v}})\coloneqq\mathbb{E}_{{\bm{z}},t,\bm{\epsilon},c}[\lVert\bm{\epsilon}-\bm{\epsilon}_{\theta}({\bm{z}}_{t},t,c({\bm{v}}))\rVert^{2}_{2}].$ (4)
Here, $\bm{\epsilon}_{\theta}$ and $c(\cdot)$ are the diffusion model and text encoder, respectively. The Euclidean gradient of this objective, ${\bm{g}}_{\text{euc}}=\nabla_{{\bm{v}}}\mathcal{L}$, is used in the RSGD update.
@@
Constant-direction prior gradient.  A useful property is that the Euclidean gradient of log-prior is a constant: $\nabla_{\bm{v}}(-\kappa\bm{\mu}^{\mathsf{T}}{\bm{v}})=-\kappa\bm{\mu}$. Practically, we just add this vector to the data gradient before projecting to the tangent space and retracting. This is analogous in spirit to decoupled weight decay (loshchilov_decoupled_2019), but adapted for the sphere with a directional prior. The update is computationally cheap (requiring no new graph operations), numerically stable, and highly interpretable: it applies a constant pull towards a semantically meaningful direction.
Selection of vMF parameters.  The vMF prior is defined by a mean direction $\bm{\mu}$ and a concentration parameter $\kappa$. The mean direction $\bm{\mu}$ is set to the normalized embedding of a corresponding class token (e.g., ‘dog’) from the pre-trained text encoder and is held constant during optimization. Since estimating $\kappa$ is non-trivial, we treat it as a hyperparameter that controls the strength of the prior. We performed a grid search and found that values in the range of 5e-5 to 2e-4 works well. Based on this, we simply fixed the value of $\kappa$ to 1e-4 for all experiments. Further discussion on the selection of prior can be found in Appendix D.2 and  D.4.
@@ 4.1 Experimental setups
Models.  Unless otherwise specified, we employed Stable Diffusion XL (SDXL) (podell_sdxl_2024) as our primary model due to its superior performance and widespread adoption in concurrent research. To demonstrate DTI’s applicability to more recent architectures, we conducted additional experiments on SANA 1.5 (xie2024sana), which employs Gemma (team2024gemma) as the text encoder and DiT (peebles2023scalable) as the image generator.
@@ 4.2 Main results
Quantitative results.  In Table 3, we quantitatively evaluate DTI along two axes: subject similarity and text–prompt fidelity. DTI consistently produces outputs that adhere closely to the prompt while maintaining high subject similarity. To isolate the role of embedding norm analyzed in Section 2.2, we rescaled TI’s learned embeddings to the in-distribution norm—specifically, the average norm of the vocabulary embeddings, matching the norm scale used in DTI. Consistent with our analysis, this simple rescaling noticeably improves text fidelity but does not fully resolve the problem, as it degrades image similarity. CrossInit achieves strong text fidelity on SDXL but fails to do so consistently on SANA, which we attribute to differences in their text encoders; SDXL uses a CLIP text encoder, while SANA employs the LLM-based encoder. Notably, DTI’s advantage over the baselines become even more pronounced as the model size increases. Overall, these results clearly demonstrates the advantage of DTI over competing baselines. Additional comparisons with further baselines on other Stable Diffusion variants are provided in Appendix D.3.
@@
Qualitative results.  Figure 3 illustrates qualitative comparisons across various prompts. DTI consistently generates images that more accurately reflect the content of the captions, while effectively preserving subject consistency. For instance, for ‘Pop-art style illustration of <cat>’, TI omits the cat while DTI renders the cat in the specified style. Similarly, in the second column, TI and CrossInit fail to incorporate all elements of the prompt, disregarding either the subject or details such as ‘music stage’ and ‘spotlight’. In contrast, DTI integrates both the subject and these details, producing a more complete output. Collectively, these examples highlight DTI’s superior compositional fidelity and subject preservation, showing its powerfulness that consistently satisfies all prompt constraints. This attributes to DTI’s stable optimization within the directional space, which facilitates improved integration of multiple prompt components. DTI’s ability to maintain subject fidelity and adhere to textual intent establishes it as a robust choice for a wide range of text-to-image generation tasks. Additional qualitative results including those of SANA can be found in Appendix D.6.
@@ 4.3 Ablation study
We performed ablation study to verify the effectiveness of components of our DTI, including the optimization space, the embedding magnitude $m$, and the concentration parameter of vMF distribution $\kappa$. The results are summarized in Table 3. To validate our choice of Riemannian SGD (RSGD), we compared it against a baseline using the AdamW optimizer. This baseline performs standard Euclidean updates and then projects the vector back onto the unit sphere after each step, which is not a true Riemannian update. The results show that RSGD substantially outperforms AdamW, highlighting the benefit of respecting the geometry of the directional manifold. Next, we found that fixing the magnitude to minimum or out-of-distribution scale has negatively affect either subject similarity or text fidelity. Setting the magnitude to an in-distribution scale yields the best results. Lastly, removing the prior (i.e., $\kappa=0$) or extremely high values of $\kappa$ hurts the performance, while moderate incorporation of prior provides the most stable results. Overall, we confirm that these ablation results validate our design choices. Further analyses are provided in Appendix D.4.

### v2

@@ 1 Introduction
Personalization in text-to-image generation involves the targeted adaptation of models to learn representations of novel, user-provided concepts (gal_image_2023; ruiz_dreambooth_2023). This process allows for the creation of customized images that faithfully render specific concepts, such as unique individuals, objects, or artistic styles, in new contexts.
Current personalization approaches fall into two paradigms: parameter fine-tuning and embedding optimization. Parameter fine-tuning methods, exemplified by DreamBooth (ruiz_dreambooth_2023), optimize entire models using a few user-provided images. While effective, these approaches are computationally expensive and require significant storage per concept. In contrast, embedding optimization methods, such as Textual Inversion (gal_image_2023), offer a more efficient alternative by optimizing only token embeddings. This approach provides substantial advantages: minimal storage per concept and seamless workflow integration. These advantages have made TI a foundational component in numerous personalization frameworks (hao_vico_2023; kumari_multi-concept_2023; tewel2023key; lee_direct_2024) and align with a broader paradigm shared with other domains, such as LLMs (lester2021power) and VLMs (alaluf2024myvlm).
Despite its utility, TI suffers from critical limitations. The fundamental challenge stems from the restrictive constraint of optimizing a single embedding vector to encapsulate highly complex and multifaceted visual concepts. This limitation leads to two key problems. First, TI struggles to maintain high fidelity to complex prompts, compromising its controllability and expressive range. Second, the extensive fine-tuning duration required to learn each new concept hinders its practical applicability for rapid, user-driven workflows. Recent works (voynov_p_2023; alaluf2023neural) have attempted to address these limitations through enriched embedding spaces, but introduce significant computational overhead that undermines TI’s efficiency advantage. Moreover, these existing methods merely treat the symptoms rather than directly addressing the underlying optimization dynamics of TI, leaving the fundamental geometric factors that govern semantic alignment in embedding-based personalization largely unclear.
To bridge this gap, this paper presents a systematic, interpretability-driven analysis of the optimization dynamics inherent to TI, with a specific focus on the geometric characteristics of the token embedding space. By carefully dissecting the learned representations, our investigation reveals that semantic information is predominantly encoded in the direction of the embedding vectors. Furthermore, we demonstrate both theoretically and empirically that the unconstrained magnitude of these embeddings is a primary source of instability; specifically, excessively high embedding norms emerge during optimization and act as a critical factor impairing image-text alignment.
@@ 1 U+2069 apples
As shown in Table 2.1, an embedding’s nearest neighbors are semantically coherent when measured by cosine similarity but not by Euclidean distance. For the token ‘apple’, its cosine-based neighbors include ‘apples’, ‘fruit’, and ‘pear’, while its Euclidean-based neighbors are often semantically unrelated tokens that merely share a similar magnitude. This indicates that an embedding’s direction is the primary carrier of semantic information. More results are provided in Appendix A.
@@ 2.2 Why large magnitudes lead to low text fidelity
Effect II: Residual updates stagnate (see Lemma 2).  The residual updates, $F_{\ell}(\operatorname{Norm}({\bm{x}}^{(\ell)}))$, are computed from normalized inputs and thus have a bounded magnitude. When this bounded update is added through the skip connection to a large vector ${\bm{x}}^{(l)}$, the relative change (i.e., turning angle of the hidden state’s direction) becomes tiny, decreasing in proportion to $1/\|{\bm{x}}^{(l)}\|$. In other words, large-norm hidden states become stuck in their direction and are difficult for subsequent layers to refine. This residual stagnation accumulates across layers, severely limiting the total directional change the initial token can undergo, as formalized in the following proposition and corollary.
@@
$\angle(\alpha{\bm{x}}^{(0)},{\bm{x}}^{(L)}(\alpha))\;\leq\;\frac{\pi}{2}\,\frac{S_{L}}{\alpha\left\lVert{\bm{x}}^{(0)}\right\rVert-S_{L}}\;\xrightarrow[\alpha\to\infty]{}\;0,$
@@
Together, these two effects explain why TI struggles with text fidelity. As a token’s magnitude increases, its ability to integrate contextual information from the prompt diminishes. The personalized token becomes so dominant that it overshadows other critical details, such as stylistic elements, background context, or additional subjects, from the generated output. To this end, this analysis highlights the need for a method that explicitly controls the magnitude of personalized tokens, which we introduce in the next section.
@@ 2.3 Empirical Validation
We evaluate whether increasing embedding magnitude makes positional information unrecoverable after the first pre-norm normalization ($\mathrm{LN}$). To validate this, we train a 2-layer MLP classifier on the frozen base text encoder to predict a token’s absolute position from the output of $\mathrm{LN}(\mathbf{e}+\mathbf{p})$, where $\mathbf{e}$ and $\mathbf{p}$ each denote token and positional embeddings. On unmodified inputs (‘Normal‘ in Figure 2), the classifier achieves $100\%$ accuracy, confirming that $\mathrm{LN}$ preserves positional information. We then scale the norm of a single token embedding by a factor $m\in\{0.5,1,2,4,8,16\}$ before applying $\mathrm{LN}$. Accuracy deteriorates rapidly once $m$ exceeds the natural scale of the encoder. Furthermore, we evaluated the classifier on TI-trained and DTI-trained personalized embeddings. TI embeddings, which have excessively large norms, collapse to near-zero positional accuracy, while DTI embeddings remain fully recoverable.
@@ 3 Method: Directional Textual Inversion
Based on the observation and analysis in the previous section that token embeddings exhibit strong directional characteristics, we introduce Directional Textual Inversion (DTI), a framework that optimizes an embedding’s direction with in-distribution norm to enhance text fidelity in personalized text-to-image generation.
@@ 3.1 Optimizing only direction on the hypersphere
Here, $\mathbb{S}^{d-1}=\{{\bm{u}}\in R^{d}:\|{\bm{u}}\|_{2}=1\}$ denotes the unit sphere. We fix the magnitude $m^{\star}$ and optimize only the direction ($\bm{v}$}). Specifically, we set $m^{\star}$ to be an in-distribution magnitude derived from the frozen vocabulary of the text encoder (e.g., the average norm). In this way, optimization focuses on semantic information in direction while avoiding out-of-distribution (OOD) norms.
Since the parameter space is the unit sphere, Euclidean updates drift off-manifold, rendering AdamW (loshchilov_decoupled_2019)–the default optimizer for TI-like methods–unsuitable. To solve this, we use Riemannian stochastic gradient descent (RSGD) (bonnabel2013rsgd) with tangent-space projection and retraction:
@@
Here, ${\bm{g}}_{\text{euc}}$ is a Euclidean space gradient, ${\bm{g}}\in T_{{\bm{v}}_{k}}\mathcal{S}^{d-1}$ is a tangent-space gradient, and $\eta>0$ is a learning rate. In practice, we further scaled $\bm{g}$} by its norm. This was inspired by Euclidean space optimizers (hinton2012rmsprop; kingma_adam_2015; loshchilov_decoupled_2019), which normalize gradients based on moving averages of squared gradients. See Algorithm 1 and Appendix C.1 for further details.
@@ 3.2 Maximum A Posteriori formulation with a directional vMF prior
To incorporate a directional prior, we formulate the optimization for the optimal direction $\bm{v}$$^{*}$ as a Maximum A Posteriori (MAP) estimation problem. Given a dataset of images $\mathcal{D}=\{{\bm{z}}_{1},\dots,{\bm{z}}_{n}\}$, the MAP estimate is found by maximizing the posterior probability:
${\bm{v}}^{*}=\arg\max_{\mathbf{v}}p(\mathbf{v}\mid\mathcal{D})=\arg\max_{\mathbf{v}}\left[\log p(\mathcal{D}\mid\mathbf{v})+\log p(\mathbf{v})\right].$ (3)
Minimizing the negative log-posterior is equivalent to minimizing a loss function composed of a data term and a prior term, $\mathcal{L}({\bm{v}})=\mathcal{L}_{\text{data}}(m^{*}{\bm{v}})+\mathcal{L}_{\text{prior}}({\bm{v}})$.
@@
$\mathcal{L}_{\text{data}}(m^{*}{\bm{v}})\coloneqq\mathbb{E}_{{\bm{z}},t,\bm{\epsilon}}[\lVert\bm{\epsilon}-\bm{\epsilon}_{\theta}({\bm{z}}_{t},t,c(m^{*}{\bm{v}}))\rVert^{2}_{2}].$ (4)
Here, $\bm{\epsilon}_{\theta}$ and $c(\cdot)$ are the diffusion model and text encoder, respectively. For notational simplicity, we write $c(m^{*}{\bm{v}})$ as shorthand for the text conditioning obtained from a sampled prompt template containing the personalized token with embedding $m^{*}{\bm{v}}$, and omit the explicit dependence on text prompt throughout the paper.
@@
Constant-direction prior gradient.  A useful property is that the Euclidean gradient of the regularization term is a constant: $\nabla_{\bm{v}}(-\kappa\bm{\mu}^{\mathsf{T}}{\bm{v}})=-\kappa\bm{\mu}$. Practically, we just add this vector to the data gradient before projecting to the tangent space and retracting. This is analogous in spirit to decoupled weight decay (loshchilov_decoupled_2019), but adapted for the sphere with a directional prior. The update is computationally cheap (requiring no new graph operations), numerically stable, and highly interpretable: it applies a constant pull towards a semantically meaningful direction.
Selection of vMF parameters.  The vMF prior is defined by a mean direction $\bm{\mu}$ and a concentration parameter $\kappa$. The mean direction $\bm{\mu}$ is set to the normalized embedding of a corresponding class token (e.g., ‘dog’) from the pre-trained text encoder and is held constant during optimization. Since estimating $\kappa$ is non-trivial, we treat it as a hyperparameter that controls the strength of the prior. We performed a grid search and found that values in the range of 5e-5 to 2e-4 work well. Based on this, we simply fixed the value of $\kappa$ to 1e-4 for all experiments. Further discussion on the selection of prior can be found in Appendix D.2 and  D.4.
@@ 4.1 Experimental setups
Models.  Unless otherwise specified, we employed Stable Diffusion XL (SDXL) (podell_sdxl_2024) as our primary model due to its superior performance and widespread adoption in recent literature. To demonstrate DTI’s applicability to more recent architectures, we conducted additional experiments on SANA 1.5 (xie2025sana), which employs Gemma (team2024gemma) as the text encoder and DiT (peebles2023scalable) as the image generator.
@@ 4.2 Main results
Quantitative results.  In Table 3, we quantitatively evaluate DTI along two axes: subject similarity and text–prompt fidelity. DTI consistently produces outputs that adhere closely to the prompt while maintaining high subject similarity. To isolate the role of embedding norm analyzed in Section 2.2, we rescaled TI’s learned embeddings to the in-distribution norm—specifically, the average norm of the vocabulary embeddings, matching the norm scale used in DTI. Consistent with our analysis, this simple rescaling noticeably improves text fidelity but does not fully resolve the problem, as it degrades image similarity. CrossInit achieves strong text fidelity on SDXL but fails to do so consistently on SANA, which we attribute to differences in their text encoders; SDXL uses a CLIP text encoder, while SANA employs the LLM-based encoder. Notably, DTI’s advantage over the baselines becomes even more pronounced as the model size increases. Overall, these results clearly demonstrate the advantage of DTI over competing baselines. Additional comparisons with further baselines on other Stable Diffusion variants are provided in Appendix D.3.
@@
Qualitative results.  Figure 3 illustrates qualitative comparisons across various prompts. DTI consistently generates images that more accurately reflect the content of the captions, while effectively preserving subject consistency. For instance, for ‘Pop-art style illustration of <cat>’, TI omits the cat while DTI renders the cat in the specified style. Similarly, in the second column, TI and CrossInit fail to incorporate all elements of the prompt, disregarding either the subject or details such as ‘music stage’ and ‘spotlight’. In contrast, DTI integrates both the subject and these details, producing a more complete output. Collectively, these examples highlight DTI’s superior compositional fidelity and subject preservation, demonstrating that it consistently satisfies all prompt constraints. We attribute this to DTI’s stable optimization within the directional space, which facilitates improved integration of multiple prompt components. DTI’s ability to maintain subject fidelity and adhere to textual intent establishes it as a robust choice for a wide range of text-to-image generation tasks. Additional qualitative results including those of SANA can be found in Appendix D.6.
@@ 4.3 Ablation study
We performed an ablation study to verify the effectiveness of components of our DTI, including the optimization space, the embedding magnitude $m$, and the concentration parameter $\kappa$ of the vMF distribution. The results are summarized in Table 3. To validate our choice of Riemannian SGD (RSGD), we compared it against a baseline using the AdamW optimizer. This baseline performs standard Euclidean updates and then projects the vector back onto the unit sphere after each step, which is not a true Riemannian update. The results show that RSGD substantially outperforms AdamW, highlighting the benefit of respecting the geometry of the directional manifold. Next, we found that fixing the magnitude to the minimum or to an out-of-distribution scale negatively affected either subject similarity or text fidelity. Setting the magnitude to an in-distribution scale yields the best results. Lastly, removing the prior (i.e., $\kappa=0$) or extremely high values of $\kappa$ hurts the performance, while moderate incorporation of the prior provides the most stable results. Overall, we confirm that these ablation results validate our design choices. Further analyses are provided in Appendix D.4.